solve the following equation using the complex number system:
x^4 + 13x^2 +36 = 0
the math book i have doesn't show how to do these types of problems: very frustrating.
the answers in the back of the book are (-2i, 2i, -3i, 3i)

Answer by Edwin McCravy(20086)

        x⁴+ 13x² + 36 = 0

Factor:

     (x² + 4)(x² + 9) = 0

x² + 4 =  0         x² + 9 = 0

    x² = -4             x² = -9
           __                  __ 
     x = ±Ö-4            x = ±Ö-9

     x = ±2i             x = ±3i 

Solutions: {-2i, 2i, -3i, 3i}

Edwin